OpenStudy (anonymous):

how do i rationalize a denominator? like -3/4 +3sq rt.6 over 1-3sq rt.6/(4)

OpenStudy (anonymous):

or tan(a-b) where a is -3/4 and b is -sq.rt6

OpenStudy (anonymous):

thanks for nothing you flutterin all suck

OpenStudy (compassionate):

I suck? I believe I was going to answer your question.

OpenStudy (anonymous):

well if you would just explain a bit id be grateful

OpenStudy (anonymous):

the last guy was very rude

OpenStudy (compassionate):

Okay, give me a second to get the right equation.

OpenStudy (compassionate):

4 Is that correct?\[/\frac{-3}{4 + 3} \sqrt{6}[Over] 1 - 3 \frac{\sqrt6}{4}\]

OpenStudy (compassionate):

I'm trying to decipher what I'm solving before I solve it.

OpenStudy (anonymous):